# homogeneous_khovanskii.bib

### v1

@@ 4. Examples
We give examples of homogeneous Khovanskii bases before we come to our main application in section 5. Although everything in this section can be computed by hand, our experiments in the computer algebra system OSCAR [DEFHJ24, Osc] were invaluable to find these examples.
@@
We give an example where Algorithm 2 is unable to find a Khovanskii basis because Algorithm 1 does not terminate. Let $K$ be a field, $B=K[x,y]$ the polynomial ring and $v:B\setminus\{0\}\to\mathbb{Z}$ be the order of divisibility by $y$ with the natural ordering on $\mathbb{Z}$. We have
$\operatorname{gr}_{v}(B)=K[x]\oplus yK[x]\oplus y^{2}K[x]\oplus\cdots=B.$
Let now $f_{1}\coloneqq x,f_{2}\coloneqq x+y,f_{3}\coloneqq y+y^{2}\in B$ and $A=K[f_{1},f_{2},f_{3}]$. The grading is unimportant for this example and we just put $\Delta\coloneqq\{0\}$. Clearly, $A=B$ and a Khovanskii basis of $A$ with respect to $v$ is given by $\{x,y\}$.
@@
$\begin{array}[]{c|c}&\operatorname{in}_{v}^{\{0\}}(\cdot)\\ \hline\cr f_{1}&x\\ f_{2}&x\\ f_{3}&y\end{array}$
@@
for all $\vartheta\in\Theta$, where $[\vartheta]$ denotes the class of $\vartheta$ in $\Delta/\Theta$.
@@
Let $\Theta\leq\Delta$ be a subgroup and let $\{f_{1},\dots,f_{k}\}$ be a $\Theta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$. If $f_{1},\dots,f_{k}$ are $\Delta$/̄homogeneous, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$.
@@
This follows directly from Theorem 2.19. Write $J_{\Delta}$, respectively $J_{\Theta}$, for the ideal $J$ coming from the respective grading. By construction, we have $J_{\Delta}\subseteq J_{\Theta}$ (any $\Delta$/̄homogeneous relations must be $\Theta$/̄homogeneous). The ideal $\operatorname{in}_{\Gamma}(I)$ in the theorem is independent of the grading by $\Delta$ or $\Theta$ and we have by assumption $\operatorname{in}_{\Gamma}(I)=J_{\Theta}$ as well as $\operatorname{in}_{\Gamma}(I)\subseteq J_{\Delta}$. Hence $\operatorname{in}_{\Gamma}(I)=J_{\Delta}$ and Theorem 2.19 gives the claim. ∎
@@
Let $f_{1}\coloneqq x^{2}+y^{2}$, $f_{2}\coloneqq x^{2}-y^{2}$ and $f_{3}\coloneqq xy$. We want to compute a $\Delta$/̄homogeneous Khovanskii basis for the subalgebra $A\coloneqq K[f_{1},f_{2},f_{3}]\leq B$ which is graded by $\Delta\coloneqq(\mathbb{Z}/2\mathbb{Z})^{2}\times\mathbb{Z}$ via $\deg_{\Delta}(f_{1})=(0,0,2)$, $\deg_{\Delta}(f_{2})=(1,0,2)$ and $\deg_{\Delta}(f_{3})=(0,1,2)$. By Proposition 2.13, Algorithm 2 is applicable because the grading refines the standard grading.
@@
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}(\cdot)&\operatorname{in}_{v}^{\Delta}(\cdot)\\ \hline\cr f_{1}&y^{2}&y^{2}g_{3}^{2}\\ f_{2}&-y^{2}&-y^{2}g_{1}g_{3}^{2}\\ f_{3}&xy&xyg_{2}g_{3}^{2}\end{array}$
@@
is given by $J=(X_{1}^{2}-X_{2}^{2})$. Hence the algorithm would check the potential new element $f_{1}^{2}-f_{2}^{2}=4x^{2}y^{2}$ with $\operatorname{in}_{v}^{\Delta}(4x^{2}y^{2})=4x^{2}y^{2}$ and the subduction algorithm rewrites this as $f_{1}^{2}-f_{2}^{2}=4f_{3}^{2}$. So, $\{f_{1},f_{2},f_{3}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$.
@@
Let us consider the same algebras $A$ and $B$ and the same valuation $v$ as in Example 4.3, but this time we take only the standard grading of $A$ (and $B$) into account and compute a $\mathbb{Z}$/̄homogeneous Khovanskii basis of $A$ with respect to $v$. We will see that the polynomials $f_{1},f_{2},f_{3}$ do not form a $\Delta$/̄homogeneous Khovanskii basis in this case.
@@
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}(\cdot)&\operatorname{in}_{v}^{\mathbb{Z}}(\cdot)\\ \hline\cr f_{1}&y^{2}&y^{2}g^{2}\\ f_{2}&-y^{2}&-y^{2}g^{2}\\ f_{3}&xy&xyg^{2}\end{array}$
@@
Evaluating the new generator gives $f_{5}\coloneqq f_{2}f_{4}+2f_{3}^{2}=2x^{4}$ with $\operatorname{in}_{v}^{\mathbb{Z}}(f_{5})=2x^{4}g^{4}$ and the subduction algorithm finds the representation $f_{5}=\frac{1}{2}f_{4}^{2}$. We conclude that $\{f_{1},\dots,f_{4}\}$ is a $\mathbb{Z}$/̄homogeneous Khovanskii basis for $A$ and $v$.
@@
For this, we adapt [KR05, Example 6.6.7]. There, Kreuzer and Robbiano show that the subalgebra generated by $x_{1}+x_{2},x_{1}x_{2},x_{1}x_{2}^{2}\in K[x_{1},x_{2}]$ has no finite SAGBI basis, no matter which monomial ordering is used on $K[x_{1},x_{2}]$. We choose a slightly different set of polynomials to produce the situation we aim for. Let $K$ be a field and let $B=K[x,y,z]$. Let $>_{1}$ be the negative lexicographical ordering on $z$, that is, induced by $1>_{1}z$, and let $>_{2}$ be a monomial ordering on the variables $x,y$ with $x>_{2}y$. Let now $>\;\coloneqq(>_{1},>_{2})$ be the block ordering, which gives a total ordering on all monomials of $B$. Note that $>_{1}$ is a local ordering (in the terminology of [GP08]) and so $>$ is out of scope of the orderings considered in [KR05] for SAGBI bases. Nevertheless, there is a valuation $v:B\setminus\{0\}\to\mathbb{Z}^{3}$ induced by $>$ as in Remark 2.7 and we can ask for Khovanskii bases with respect to $v$. We again have $\operatorname{gr}_{v}(B)\cong B$. Let finally $f_{1}\coloneqq x+y+z$, $f_{2}\coloneqq xy$ and $f_{3}\coloneqq xy^{2}$ and put $A\coloneqq K[f_{1},f_{2},f_{3}]\leq B$. The polynomials $f_{1}$, $f_{2}$ and $f_{3}$ are algebraically independent and we may endow $A$ with a grading by $\Delta\coloneqq\mathbb{Z}^{3}$ by putting $\deg_{\Delta}(f_{i})=e_{i}$, where $e_{i}\in\mathbb{Z}^{3}$ denotes the element with entry 1 in position $i$ and 0 everywhere else. Then we have the group ring $K\Delta=K[g_{1},g_{2},g_{3}]$ and one computes:
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}(\cdot)&\operatorname{in}_{v}^{\Delta}(\cdot)\\ \hline\cr f_{1}&x&xg_{1}\\ f_{2}&xy&xyg_{2}\\ f_{3}&xy^{2}&xy^{2}g_{3}\end{array}$
Hence there is no relation between the $\operatorname{in}_{v}^{\Delta}(f_{i})$, $i=1,2,3$, and a $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$ is given by $\{f_{1},f_{2},f_{3}\}$ (Algorithm 2 is applicable by Proposition 2.13).
We now consider the standard grading by $\mathbb{Z}$ on $A$ and prove in the following that there is no finite $\mathbb{Z}$/̄homogeneous Khovanskii basis of $A$. Let
@@
be the canonical projection. Then $\rho(A)$ is the algebra considered in [KR05, Example 6.6.7]. Further, by construction of $>$, we have $\operatorname{LM}_{>_{2}}(\rho(f))=\rho(\operatorname{LM}_{>}(f))$. Hence $\rho$ induces a surjection $\operatorname{in}_{v}^{\mathbb{Z}}(A)\to\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ where $v_{2}$ is the valuation on $K[x,y]$ induced by $>_{2}$. Kreuzer and Robbiano prove that there is an isomorphism $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))\cong K[x,xy,xy^{2},\dots]$ and conclude that $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ is not a finitely generated algebra. Precisely, the monomial $xy^{d}$ is not in the algebra $K[x,xy,\dots,xy^{d-1}]$ for every $d\geq 1$. The same holds true for $\operatorname{in}_{v}^{\mathbb{Z}}(\rho(A))$ because the monomials in $\operatorname{in}_{v}^{\mathbb{Z}}(\rho(A))$ which are not in $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ are all divisible by $z$. In conclusion, there is no finite $\mathbb{Z}$/̄homogeneous Khovanskii basis for $A$ with respect to $v$.
@@ 5.1. Quotient singularities
Throughout this section, let $V$ be a finite/dimensional vector space over $\mathbb{C}$ and let $G\leq\operatorname{SL}(V)$ be a finite group. We write $V/G\coloneqq\operatorname{Spec}\mathbb{C}[V]^{G}$ for the linear quotient of $V$ by $G$, where $\mathbb{C}[V]^{G}$ is the invariant ring of $G$. By the classical theorem of Chevalley–Serre–Shephard–Todd [Ser68, Théorème 1’], the variety $V/G$ is singular and we hence refer to it as a quotient singularity.
We are interested in the birational geometry of $V/G$. In the following, we give a terse summary of the steps that lead us to a certain Cox ring; we refer the reader to the given references for more details. By deep results of the minimal model programme [BCHM10], there exists a minimal model, or more precisely a $\mathbb{Q}$/̄factorial terminalization, $X\to V/G$ of $V/G$, see [Sch23, Theorem 2.1.15] for details of how this follows from [BCHM10, Corollary 1.4.3]. Further, this minimal model turns out to be a (relative) Mori dream space [Gra19, Theorem 3.4.10]. Algebraically, this means that the Cox ring $\mathcal{R}(X)$ is a finitely generated $\mathbb{C}$/̄algebra. This makes it accessible from an algorithmic point of view.
@@
To describe the setting, we start with the Cox ring $\mathcal{R}(V/G)$ of $V/G$. By a theorem of Arzhantsev–Gaǐfullin [AG10, Theorem 3.1], we have $\mathcal{R}(V/G)\cong\mathbb{C}[V]^{[G,G]}$, where $[G,G]$ denotes the commutator subgroup of $G$. The ring $\mathcal{R}(V/G)$ is graded by the class group $\operatorname{Cl}(V/G)$ and we have $\operatorname{Cl}(V/G)=\operatorname{Hom}(G,\mathbb{C}^{\times})$ by [Ben93, Theorem 3.9.2]. Let $\operatorname{Ab}(G)\coloneqq G/[G,G]$ be the abelianization of $G$ and write $\operatorname{Ab}(G)^{\vee}$ for the group of characters of $\operatorname{Ab}(G)$. By elementary character theory, there is an isomorphism $\operatorname{Hom}(G,\mathbb{C}^{\times})\cong\operatorname{Ab}(G)^{\vee}$. The isomorphism $\mathcal{R}(V/G)\cong\mathbb{C}[V]^{[G,G]}$ is $\operatorname{Ab}(G)^{\vee}$/̄graded, where the grading on $\mathbb{C}[V]^{[G,G]}$ is coming from the action of $\operatorname{Ab}(G)$, that is,
@@
Let $f_{1},\dots,f_{k}\in\mathbb{C}[V]^{[G,G]}$ be $\operatorname{Ab}(G)^{\vee}$/̄homogeneous generators. Then the Cox ring $\mathcal{R}(X)$ is generated by the preimages of
@@
under $\Theta$ if and only if $\{f_{1},\dots,f_{k}\}$ is a $\operatorname{Ab}(G)^{\vee}$/̄homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$.

### v2

@@ 4. Examples
We give further examples of homogeneous Khovanskii bases before we come to our main application in section 5. Although everything in this section can be computed by hand, our experiments in the computer algebra system OSCAR [Osc25, Dec+25] were invaluable to find these examples.
@@
We give an example where Algorithm 2 is unable to find a Khovanskii basis because Algorithm 1 does not terminate. Let $K$ be a field, $B=K[x,y]$ the polynomial ring and $v:B\setminus\{0\}\to\mathbb{Z}$ be the order of divisibility by $y$ with the natural ordering on $\mathbb{Z}$. We have $\operatorname{gr}_{v}(B)=B.$ Let $f_{1}\coloneqq x,f_{2}\coloneqq x+y,f_{3}\coloneqq y+y^{2}\in B$ and $A=K[f_{1},f_{2},f_{3}]$. The grading is unimportant for this example and we just put $\Delta\coloneqq\{0\}$. Clearly, $A=B$ and a Khovanskii basis of $A$ with respect to $v$ is given by $\{x,y\}$.
@@
$\begin{array}[]{c|c}&\operatorname{in}_{v}^{\{0\}}\\ \hline\cr f_{1}&x\\ f_{2}&x\\ f_{3}&y\end{array}$
@@
for all $\vartheta\in\Theta$, where $\vartheta]$ denotes the class of $\vartheta$ in $\Delta/\Theta$.
@@
Let $\Theta\leq\Delta$ be a subgroup and let $\{f_{1},\dots,f_{k}\}$ be a $\Theta$-homogeneous Khovanskii basis of $A$ with respect to $v$. If $f_{1},\dots,f_{k}$ are $\Delta$-homogeneous, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$.
@@
This follows directly from Theorem 2.21. Write $J_{\Delta}$, respectively $J_{\Theta}$, for the ideal $J$ coming from the respective grading. By construction, we have $J_{\Delta}\subseteq J_{\Theta}$ (any $\Delta$-homogeneous relations must be $\Theta$-homogeneous). The ideal $\operatorname{in}_{\Gamma}(I)$ in the theorem is independent of the grading by $\Delta$ or $\Theta$ and we have by assumption $\operatorname{in}_{\Gamma}(I)=J_{\Theta}$ as well as $\operatorname{in}_{\Gamma}(I)\subseteq J_{\Delta}$. Hence $\operatorname{in}_{\Gamma}(I)=J_{\Delta}$ and Theorem 2.21 gives the claim. ∎
@@
Let $f_{1}\coloneqq x^{2}+y^{2}$, $f_{2}\coloneqq x^{2}-y^{2}$ and $f_{3}\coloneqq xy$. We want to compute a $\Delta$-homogeneous Khovanskii basis for the subalgebra $A\coloneqq K[f_{1},f_{2},f_{3}]\leq B$ which is graded by $\Delta\coloneqq(\mathbb{Z}/2\mathbb{Z})^{2}\times\mathbb{Z}$ via $\deg_{\Delta}(f_{1})=(0,0,2)$, $\deg_{\Delta}(f_{2})=(1,0,2)$ and $\deg_{\Delta}(f_{3})=(0,1,2)$. By Proposition 2.15, Algorithm 2 is applicable because the grading refines the standard grading.
@@
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}&\operatorname{in}_{v}^{\Delta}\\ \hline\cr f_{1}&y^{2}&y^{2}g_{3}^{2}\\ f_{2}&-y^{2}&-y^{2}g_{1}g_{3}^{2}\\ f_{3}&xy&xyg_{2}g_{3}^{2}\end{array}$
@@
is given by $J=(X_{1}^{2}-X_{2}^{2})$. Hence the algorithm would check the potential new element $f_{1}^{2}-f_{2}^{2}=4x^{2}y^{2}$ with $\operatorname{in}_{v}^{\Delta}(4x^{2}y^{2})=4x^{2}y^{2}$ and the subduction algorithm rewrites this as $f_{1}^{2}-f_{2}^{2}=4f_{3}^{2}$. So, $\{f_{1},f_{2},f_{3}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$.
@@
Let us consider the same algebras $A$ and $B$ and the same valuation $v$ as in Example 4.3, but this time we take only the standard grading of $A$ (and $B$) into account and compute a $\mathbb{Z}$-homogeneous Khovanskii basis of $A$ with respect to $v$. We will see that the polynomials $f_{1},f_{2},f_{3}$ do not form a $\Delta$-homogeneous Khovanskii basis in this case.
@@
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}&\operatorname{in}_{v}^{\mathbb{Z}}\\ \hline\cr f_{1}&y^{2}&y^{2}g^{2}\\ f_{2}&-y^{2}&-y^{2}g^{2}\\ f_{3}&xy&xyg^{2}\end{array}$
@@
Evaluating the new generator gives $f_{5}\coloneqq f_{2}f_{4}+2f_{3}^{2}=2x^{4}$ with $\operatorname{in}_{v}^{\mathbb{Z}}(f_{5})=2x^{4}g^{4}$ and the subduction algorithm finds the representation $f_{5}=\frac{1}{2}f_{4}^{2}$. We conclude that $\{f_{1},\dots,f_{4}\}$ is a $\mathbb{Z}$-homogeneous Khovanskii basis for $A$ and $v$.
@@
For this, we adapt [KR05, Example 6.6.7]. There, Kreuzer and Robbiano show that the subalgebra generated by $x_{1}+x_{2},x_{1}x_{2},x_{1}x_{2}^{2}\in K[x_{1},x_{2}]$ has no finite SAGBI basis, no matter which monomial ordering is used on $K[x_{1},x_{2}]$. We choose a slightly different set of polynomials to produce the situation we aim for. Let $K$ be a field and let $B=K[x,y,z]$. Let $_{1}$ be the negative lexicographical ordering on $z$, that is, induced by $1>_{1}z$, and let $_{2}$ be a monomial ordering on the variables $x,y$ with $x>_{2}y$. Let now $\coloneqq(>_{1},>_{2})$ be the block ordering, which gives a total ordering on all monomials of $B$. Note that $_{1}$ is a local ordering (in the terminology of [GP08]) and so >> is out of scope of the orderings considered in [KR05] for SAGBI bases. Nevertheless, there is a valuation $v:B\setminus\{0\}\to\mathbb{Z}^{3}$ induced by >> as in Remark 2.8 and we can ask for Khovanskii bases with respect to $v$. We again have $\operatorname{gr}_{v}(B)\cong B$. Let finally $f_{1}\coloneqq x+y+z$, $f_{2}\coloneqq xy$ and $f_{3}\coloneqq xy^{2}$ and put $A\coloneqq K[f_{1},f_{2},f_{3}]\leq B$. The polynomials $f_{1}$, $f_{2}$ and $f_{3}$ are algebraically independent and we may endow $A$ with a grading by $\Delta\coloneqq\mathbb{Z}^{3}$ by putting $\deg_{\Delta}(f_{i})=e_{i}$, where $e_{i}\in\mathbb{Z}^{3}$ denotes the element with entry 1 in position $i$ and 0 everywhere else. Then we have the group ring $K\Delta=K[g_{1},g_{2},g_{3}]$ and one computes:
$\begin{array}[]{c|c|c}&\operatorname{in}_{v}&\operatorname{in}_{v}^{\Delta}\\ \hline\cr f_{1}&x&xg_{1}\\ f_{2}&xy&xyg_{2}\\ f_{3}&xy^{2}&xy^{2}g_{3}\end{array}$
Hence there is no relation between the $\operatorname{in}_{v}^{\Delta}(f_{i})$, $i=1,2,3$, and a $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$ is given by $\{f_{1},f_{2},f_{3}\}$ (Algorithm 2 is applicable by Proposition 2.15).
We now consider the standard grading by $\mathbb{Z}$ on $A$ and prove in the following that there is no finite $\mathbb{Z}$-homogeneous Khovanskii basis of $A$. Let
@@
be the canonical projection. Then $\rho(A)$ is the algebra considered in [KR05, Example 6.6.7]. Further, by construction of >>, we have $\operatorname{LM}_{>_{2}}(\rho(f))=\rho(\operatorname{LM}_{>}(f))$. Hence $\rho$ induces a surjection $\operatorname{in}_{v}^{\mathbb{Z}}(A)\to\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ where $v_{2}$ is the valuation on $K[x,y]$ induced by $_{2}$. Kreuzer and Robbiano prove that there is an isomorphism $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))\cong K[x,xy,xy^{2},\dots]$ and conclude that $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ is not a finitely generated algebra. Precisely, the monomial $xy^{d}$ is not in the algebra $K[x,xy,\dots,xy^{d-1}]$ for every $d\geq 1$. The same holds true for $\operatorname{in}_{v}^{\mathbb{Z}}(\rho(A))$ because the monomials in $\operatorname{in}_{v}^{\mathbb{Z}}(\rho(A))$ which are not in $\operatorname{in}_{v_{2}}^{\mathbb{Z}}(\rho(A))$ are all divisible by $z$. In conclusion, there is no finite $\mathbb{Z}$-homogeneous Khovanskii basis for $A$ with respect to $v$.
@@ 5.1. Quotient singularities
Throughout this section, let $V$ be a finite-dimensional vector space over $\mathbb{C}$ and let $G\leq\operatorname{SL}(V)$ be a finite group. We write $V/G\coloneqq\operatorname{Spec}\mathbb{C}[V]^{G}$ for the linear quotient of $V$ by $G$, where $\mathbb{C}[V]^{G}$ is the invariant ring of $G$. By the classical theorem of Chevalley–Serre–Shephard–Todd [Ser68, Théorème 1’], the variety $V/G$ is singular and we hence refer to it as a quotient singularity.
We are interested in the birational geometry of $V/G$. In the following, we give a terse summary of the steps that lead us to a certain Cox ring; we refer the reader to the given references for more details. By deep results of the minimal model programme [BCHM10], there exists a minimal model, or more precisely a $\mathbb{Q}$-factorial terminalization, $X\to V/G$ of $V/G$, see [Sch23, Theorem 2.1.15] for details of how this follows from [BCHM10, Corollary 1.4.3]. Further, this minimal model turns out to be a (relative) Mori dream space [Gra19, Theorem 3.4.10]. Algebraically, this means that the Cox ring $\mathcal{R}(X)$ is a finitely generated $\mathbb{C}$-algebra. This makes it accessible from an algorithmic point of view.
@@
To describe the setting, we start with the Cox ring $\mathcal{R}(V/G)$ of $V/G$. By a theorem of Arzhantsev–Gaǐfullin [AG10, Theorem 3.1], we have $\mathcal{R}(V/G)\cong\mathbb{C}[V]^{[G,G]}$, where $[G,G]$ denotes the commutator subgroup of $G$. The ring $\mathcal{R}(V/G)$ is graded by the class group $\operatorname{Cl}(V/G)$ and we have $\operatorname{Cl}(V/G)=\operatorname{Hom}(G,\mathbb{C}^{\times})$ by [Ben93, Theorem 3.9.2]. Let $\operatorname{Ab}(G)\coloneqq G/[G,G]$ be the abelianization of $G$ and write $\operatorname{Ab}(G)^{\vee}$ for the group of characters of $\operatorname{Ab}(G)$. By elementary character theory, there is an isomorphism $\operatorname{Hom}(G,\mathbb{C}^{\times})\cong\operatorname{Ab}(G)^{\vee}$. The isomorphism $\mathcal{R}(V/G)\cong\mathbb{C}[V]^{[G,G]}$ is $\operatorname{Ab}(G)^{\vee}$-graded, where the grading on $\mathbb{C}[V]^{[G,G]}$ is coming from the action of $\operatorname{Ab}(G)$, that is,
@@
Let $f_{1},\dots,f_{k}\in\mathbb{C}[V]^{[G,G]}$ be $\operatorname{Ab}(G)^{\vee}$-homogeneous generators. Then the Cox ring $\mathcal{R}(X)$ is generated by the preimages of
@@
under $\Theta$ if and only if $\{f_{1},\dots,f_{k}\}$ is a $\operatorname{Ab}(G)^{\vee}$-homogeneous MUVAK basis of $\mathbb{C}[V]^{[G,G]}$ with respect to $v_{1},\dots,v_{m}$.